1. If you had 6.50 moles of ferric hydroxide how many grams of ferric sulfate will be produced?


2. If 25.7 grams of water was produced what is the amount of sulfuric acid needed to produce the

water?



1
Expert's answer
2021-06-03T05:27:31-0400

a)4FeSO4 + 3H2O-->> 2Fe2(OH)3 + 4SO2 + SO3

Mole ratio = 4FeSO4 = 2Fe2O3

Molar Mass of FeSO4 = 151.908

= 6.50 × 2 × 151.908

= 1974.804g

b) Zn + H2SO4 → ZnSO4 + H2O

Mole ratio = 1:1

Molar Mass of H2O = 18.01528

= 25.7/18.01528

= 1.43 moles

Molar Mass of H2SO4 = 98.079

= 1.43 × 98.079

= 139.92g